OpenStudy (anonymous):

lim x->0 sinx/(x+tanx)

OpenStudy (freckles):

hint : \[\frac{\sin(x)}{x+\frac{\sin(x)}{\cos(x)}} =\frac{\frac{\sin(x)}{x}}{\frac{x}{x}+\frac{1}{\cos(x)} \frac{\sin(x)}{x}}\]

OpenStudy (anonymous):

1/2 Math is amazing. Thanks.

OpenStudy (freckles):

yep

OpenStudy (anonymous):

New way of approaching: when you see a sinx with limx->0, divide it by x and see what happens.

OpenStudy (freckles):

lol yep

OpenStudy (freckles):

i guess that was something you just enter into your notes for the future some time

OpenStudy (freckles):

do you know how to evaluate this limit: \[\lim_{x \rightarrow 0}\frac{\sin(5x)}{x}\]